# Dual-Flow: Transferable Multi-Target, Instance-Agnostic Attacks via In-the-wild Cascading Flow Optimization

### v1

@@ Appendix E More Analyses
Table 6: Comparison of Dual-Flow and Dual-Flow-$L_{2}$. The surrogate model is Res-152.
@@
Appendix F More Visualization
Figure 6: Visualization results of different input images targeting various classes. For each text prompt of the target class, the left column displays the adversarial examples generated before clipping, the middle column shows the adversarial examples after clipping, and the right column presents the corresponding adversarial perturbations, which represent the differences between the clipped adversarial examples and the original images. Note that the perturbations are scaled to a range between 0 and 1. The surrogate model used is Inc-v3.

### v3

@@ Appendix H More Ablation Studies
Table 14: Comparison of Dual-Flow and Dual-Flow-$L_{2}$. The source model is Res-152.
@@ Appendix I More Analysis
Appendix J More Visualization
Figure 6: Visualization results of different input images targeting various classes. For each text prompt of the target class, the left column displays the adversarial examples generated before clipping, the middle column shows the adversarial examples after clipping, and the right column presents the corresponding adversarial perturbations, which represent the differences between the clipped adversarial examples and the original images. Note that the perturbations are scaled to a range between 0 and 1. The source model used is Inc-v3.
Appendix K Limitations
While our proposed Dual-Flow framework demonstrates strong adversarial attack performance and transferability, several limitations remain. First, the training process involves additional computational overhead compared to some simpler attack methods, which may affect scalability to extremely large datasets or models. Second, although we observe improved semantic consistency in generated perturbations, there is still room to enhance the interpretability and controllability of adversarial patterns in more complex scenarios. Lastly, our evaluations focus primarily on standard image classification benchmarks; extending the approach to other tasks or modalities warrants further investigation.
Appendix L Statistical Significance
To rigorously evaluate the reliability of our attack success rates, we partitioned the test dataset into 5 disjoint subsets, each containing 200 images. We computed the attack success rate for each subset independently and derived the overall 95% confidence intervals (CIs) for our method. These confidence intervals provide a quantitative measure of variability and statistical uncertainty. We further compared our results against the baseline method by examining the overlap of their respective confidence intervals. As shown in Table 16, the non-overlapping intervals observed in our experiments indicate that the improvement in attack success rate achieved by our method is statistically significant with high confidence.
Table 16: Attack success rates with 95% confidence intervals over 5 data splits, compared to the baseline method. The source model is Res-152.
Method Inc-v3 Inc-v4 Inc-Res-v2 DN-121 GoogleNet VGG-16
[…]
Appendix M Societal Impacts
Our work contributes to a deeper understanding of adversarial vulnerabilities in deep learning models, which is essential for improving model robustness and security. By developing more effective attack methods, we provide valuable tools for evaluating and strengthening defenses against malicious exploitation. However, as with any adversarial technique, there is a potential risk of misuse in compromising AI systems. We advocate for responsible use of such methods strictly within research and security auditing contexts, and encourage the community to develop corresponding mitigation strategies to safeguard AI applications.